OpenStudy (goformit100):

Mr. John was x years old in year x^2. When was John born if he died in 1871.

OpenStudy (goformit100):

@terenzreignz @TessaMarie08181997 @mikaela19900630

OpenStudy (compassionate):

x^2 - x = 1871, right?

OpenStudy (goformit100):

Ok Thank you madam As its Quadratic equation I solved it using root method

OpenStudy (compassionate):

I appreciate the medals and all , but I'm not sure if it's the right answer you're looking for. I know if you take the year x^2 and subtract his age, you'll get his date of birth. Then given the time he died, you can just do the math.

OpenStudy (compassionate):

It's like, if I'm 15 and the year is 2011, then I subtract 2011 - 15 and my date of birth = 1996. Then if I died in 2012, then I died when I was 16.

OpenStudy (goformit100):

ok Madam

OpenStudy (anonymous):

x^2-x=18(something, probably) 42^2=1740 43^2=1849 (seems likely) makes him 32 years old (possibly 31, if he died earlier in the year than he was born)